### v1

# Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach

Abstract
With the emergence of 6G, mobile networks are becoming increasingly heterogeneous and dynamic, necessitating advanced automation for efficient management. Intent-Driven Networks (IDNs) address this by translating high-level intents into optimization policies. Large Language Models (LLMs) can enhance this process by understanding complex human instructions to enable adaptive, intelligent automation. Given the rapid advancements in Generative AI (GenAI), a comprehensive survey of LLM-based IDN architectures in disaggregated Radio Access Network (RAN) environments is both timely and critical. This article provides such a survey, along with a case study on a hierarchical learning-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution. Unlike most existing approaches that apply GenAI in the form of LLMs for intent processing only, we propose a hierarchical framework that introduces GenAI across all three stages of IDN. To demonstrate the effectiveness of the proposed IDN management architecture, we present a case study based on the latest GenAI architecture named Mamba. The case study shows how the proposed GenAI-driven architecture enhances network performance through intelligent automation, surpassing the performance of the conventional IDN architectures.
Index Terms:
Intent processing, Intent validation, Intent execution, Generative AI, Hierarchical learning
I Introduction
Sixth-Generation (6G) networks are anticipated to support a diverse set of user requirements and have more complex deployments [1]. Traditional network management methods, which are based on manual configurations and human expertise, will face challenges when optimizing and operating such complex networks. Furthermore, manual configurations are costly, prone to errors and not scalable [2]. Intent-Driven Network (IDN) management is emerging as a solution to these challenges [1]. An intent defines desired outcomes by specifying target metrics without detailing implementation steps to ensure both human comprehension and machine interpretability [4]. Advances in Natural Language Processing (NLP) help operators express these intents more effectively. This reduces manual work, lowers errors, and speeds up configuration changes and service deployments [4, 5].
Disaggregated Radio Access Networks (RANs), like Open RAN with its layered structure, are well-suited for IDN management. High-level controllers can interpret operator goals and check if the intent is achievable based on real-time conditions. Low-level controllers or functions then create the best network policies to meet these goals. This approach improves flexibility, scalability, and decision-making efficiency.
On the other hand, Generative Artificial Intelligence (GenAI) has led to the rapid advancement of Large Language Models (LLMs) and revolutionized the field of NLP. As they have an outstanding ability to understand complex human instructions in natural language, IDN management is leaning towards LLMs instead of using fixed Service Level Agreement (SLA)-based approaches. For example, Dzeparoska et al. have introduced a pipeline that utilizes an LLM to process intents into structured and policy-based abstractions, which are next linked with Application Programming Interfaces (APIs) for execution [6]. Also, in our previous work we have introduced an IDN management framework that uses LLM few-shot learning for intent processing from a human network operator [4]. While these works use LLMs for intent processing, it can be highly beneficial to introduce GenAI not only as an intent processor via LLM but also we can take IDN management to the next level by introducing GenAI solutions in the other two stages namely intent validation and intent execution.
Figure 1: Three-step methodology for intent-driven network management.
With the inspiration for introducing efficient GenAI-based IDN management schemes via hierarchical disaggregated RAN environments, this work first reviews the state-of-the-art approaches that use GenAI for IDN management. Next, we introduce a three-step methodology for IDN management, encompassing intent processing, validation, and execution (see Fig. 1 for details). After that, we present how these three key steps can be implemented using a hierarchical RAN architecture with the help of multiple GenAI algorithms. Lastly, we present a case study that shows the implementation of these steps. In particular, we process intents in a memory-efficient manner where an LLM is fine-tuned with a custom dataset using Quantized Low Rank Adaptation (QLoRA) [7]. For the intent validation, a transformer-based time series predictor has been used in the case study which predicts crucial network parameters in future time slots and based on the predictions validates intents. Lastly, for intent execution, we generate network policies using Mamba, a GenAI architecture based on state space models for its demonstrated effectiveness compared to the transformer architecture in terms of memory and computation efficiency [8].
The core contributions of this paper are as follows: it presents a precise, yet comprehensive survey of LLM-based IDN architectures, highlighting current limitations and identifying opportunities for integrating broader GenAI techniques across the stages of intent processing, validation, and execution. Furthermore, it introduces a three-step methodology for IDN management that encompasses these stages, integrated within a hierarchical RAN architecture. To validate the proposed framework, a case study demonstrating the utility of the proposed framework using decision Mamba has been presented for the first time in the literature.
II GenAI-based Intent-driven Network Management Schemes
Intent-based networking connects end-user service requests with network operations. Nijah et al. states that NLP-based methods offer the most flexibility to enable intent expression in natural language [2]. LLMs have revolutionized NLP and they surpass traditional methods in flexibility and accuracy. Their ability to perform zero-shot, few-shot, and fine-tuned learning makes them state-of-the-art for tasks like intent recognition and processing. Existing research on IDN management schemes vastly uses LLMs to understand intents from the operators. While a few GenAI paradigms explore techniques like diffusion models and neuro-symbolic reasoning, the majority of current research in IDN management focuses on LLMs. Therefore, in this section, we review the LLM-centric approaches for IDN management in the telecommunication domain. Subsequently, we explore the potential application of LLM techniques or GenAI algorithms which to date, have not been widely explored in the literature. We also summarize the LLM-centric IDN management approaches, their main features, challenges, and applications in Table I.
TABLE I: LLM-Based IDN Management Approaches: Features, Challenges, and Applications
Category Main Features Challenges Applications
LLM-assisted intent processing and management for Fifth-Generation (5G) core networks [9, 10] Supports multiple intent transmission to the model. Enhances network orchestration with semantic routing. Prevents LLM hallucination. Includes a RAG module. Reliance on LLMs which are not inherently designed for network control and automation. Distinguishing between similar or overlapping intents. Conversion of operator-defined intents into slice deployment policies. Mapping intents to edge-based service deployments.
LLM-Driven multi-agent and negotiation-based intent management for 6G networks [11, 5] Employs MARL and LLM-driven negotiation. Resolves conflicting intents in multi-stakeholder settings. Cross-domain conflict resolution. Fairness in resource allocation. Scalability concerns. 6G network slicing. Dynamic spectrum allocation. Autonomous resource management.
End-to-end AI-based intent-driven network automation [1, 2, 4] Converts high-level intents into actionable network policies. Verifies if an intent is feasible given the network’s current or predicted state. End-to-end automation using AI. Computational complexity of processing and validating multiple intents in real-time increases exponentially. Zero-touch network configuration. Predictive maintenance and fault recovery.
Generative AI and multimodal intent-based network management [12] Uses multimodal generative AI for intent translation. LLMs are fine-tuned with few-shot learning to adapt to networking use cases. High dependency on predefined templates. Expansion to include multimodal inputs (e.g., images, topology files). Automation of the deployment and management of network slices.
II-A State-of-the-art LLM-Based Approaches for IDN Management
II-A1 LLM-Assisted Intent Processing and Management for 5G Core Networks
Two recent studies explore LLM-assisted approaches on intent-based management in 5G core networks. Manias et al. in their work [9] explore LLM-driven intent extraction as a key enabler of next-generation zero-touch network service management. A customized LLM model is designed to interpret and translate user intents into actionable network policies. By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks, such as deployment, modification, performance assurance, and feasibility checking.
Semantic routing is introduced to refine LLM-assisted intent-based networking in [10]. Traditional LLM-driven approaches face issues such as hallucination, scalability limitations, and accuracy degradation when handling complex network intents. To address these challenges, the proposed framework integrates a semantic router, which ensures deterministic decision-making by routing extracted intents through predefined pathways rather than relying solely on LLM-generated responses. This hybrid architecture improves reliability, reduces latency, and enhances accuracy in 5G core management and orchestration.
II-A2 LLM-Driven Multi-Agent and Negotiation-Based Intent Management
The framework proposed in [11] is a collaborative, multi-agent system for managing shared network resources in 6G. This system deploys LLM-based agents to represent different business entities. Each of these entities negotiates service-level objectives such as throughput, cost efficiency, and energy savings. The framework acts as a central mediator to resolve conflicts by utilizing LLMs alongside optimization techniques and real-time network observability.
Another study [5] introduces a comprehensive LLM-driven intent life cycle management system designed to handle all stages of intent processing, from decomposition and translation to negotiation, activation, and assurance. By shifting away from rigid JSON/YAML intent definitions, the system proposed in [5] enables natural language-driven network orchestration. Users can specify operational goals without requiring deep technical expertise.
II-A3 End-to-End AI-Based Intent-Driven Network Automation
Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), Multi-Agent RL (MARL), and Hierarchical RL (HRL) to realize fully automated network management.
While processing intents is crucial, ensuring their feasibility and impact is equally important [4]. A recent study introduces a transformer-based time series predictor to validate intents before execution. This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations (e.g., increasing energy efficiency, improving throughput) do not negatively affect service quality. Once an intent is validated, an HRL framework selects and triggers appropriate network optimization applications such as beamforming, traffic steering, and power control. An attention-based HRL model filters out suboptimal actions to reduce computational overhead while maximizing efficiency.
Another breakthrough in end-to-end AI-driven network automation is the integration of multi-agent learning frameworks [1], where AI-driven network agents negotiate and resolve conflicting intents in real-time. This is particularly essential in multi-tenant 6G networks, where various stakeholders (e.g., operators, service providers, enterprise users) have competing resource demands. The study conducted in [1] proposes a multi-domain orchestration model, where LLM-driven agents representing RAN, core, and transport domains collaborate to optimize resource allocation and maintain Quality-of-Service (QoS).
II-A4 Generative AI and Multimodal Intent-Based Network Management
A multimodal intent recognition framework presented in [12] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators to specify intents through text prompts, graphical designs, or configuration files. AI dynamically refines intents to ensure accuracy while aligning with service-level objectives.
A key application of this approach is network-as-a-service orchestration, where LLM-powered agents automate network slice management. The framework enables zero-touch deployment to map business requirements to network slice templates.
II-B Potential GenAI Paradigms for Enhancing IDN Management
The key limitations of the LLM-based IDN management schemes that we discussed in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements. To address these issues, we recommend some possible solutions from the GenAI perspective.
$\bullet$ Retrieval Augmented Generation (RAG) and separate intent refinement/ validation scheme: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination. Furthermore, a multi-step intent refinement process can be introduced where a second model (e.g., a smaller LLM or rule-based checker) cross-checks intent translations before execution.
$\bullet$ Reinforcement Learning with Human Feedback (RLHF): Since re-training is not always an option due to computational resource scarcity, RLHF can be a powerful technique for improving LLM-based IDN management to address the issue of real-time adaptation and feedback loops. We can define a reward function based on network Key Performance Indicators (KPIs) (e.g., latency, throughput, energy efficiency). Then, a feedback loop can be implemented where operators and simulated environments score the LLM ’s intent outputs. Lastly, RL algorithms can be used to refine LLM-generated responses.
$\bullet$ LLM quantization and compression: Computation overhead caused by LLM deployments can be handled by implementing 4-bit quantization techniques (QLoRA, Generative Pre-trained Transformer-Quantization (GPTQ)) to reduce LLM model size while maintaining high accuracy.
III Three-step Methodology for Intent-driven Network Management
In this section of the paper, we present a three-step framework to introduce IDN management in modern-day communication systems. In particular, we propose a three-fold strategy consisting of intent processing, intent validation, and intent execution via network optimization policy generation.
Figure 2: Integration of the proposed IDN management framework with hierarchical RAN architecture.
$\bullet$ Intent processing: Intent processing is the phase where high-level, human-expressed network intents (e.g., “Minimize latency for video traffic”) are interpreted, extracted, and translated into structured, machine-executable commands using LLMs, multimodal AI, or rule-based systems. It is the first step of the IDN management and is highly important because if intent interpretation is not done correctly, it can lead to performance degradation.
$\bullet$ Intent validation: Intent validation is the process of assessing whether the current network conditions and capacity can accommodate a given intent or a set of intents. This involves verifying whether the available resources, traffic conditions, and operational constraints allow for the successful execution of the desired network optimization actions. Intent validation ensures that a processed intent is feasible, conflict-free, and optimal before execution.
$\bullet$ Intent execution: Intent execution is the final stage of IDN management, where validated intents are transformed into real-time network actions. This phase involves invoking appropriate network optimization policies to dynamically configure resources, and ensure that the desired service-level objectives are met efficiently. Intent execution can be realized through network function orchestration, AI-driven decision-making, or Reinforcement Learning (RL)-based adaptation. There are multiple ways to execute intents depending on the network architecture and management framework. The system can trigger predefined network applications such as traffic steering, beamforming, power control, and slicing management for intent execution. Also, LLMs can generate and refine network policies dynamically. Another approach for multi-operator environments can be to utilize RL-based agents that can negotiate optimal execution strategies to ensure efficient network optimization.
To represent this three-step methodology, we presented Fig.1 in the Introduction, which illustrates the step-by-step process of intent-driven network management. The process begins with intent processing, where network intents can be expressed in natural language, structured formats (JSON, XML), or multi-modal data such as topology diagrams and configuration files. To process the intents, different methodologies can be applied, including LLMs for natural language understanding, rule-based approaches using predefined logic, graph-based methods for structured representation of dependencies, or finite state machines to model intent workflows as a sequence of state transitions.
Once an intent is processed, it moves to intent validation which can be carried out using rule-based validation (checking against static policies), constraint-based validation (ensuring resource availability and compliance), predictive validation (using AI/ML models to anticipate network conditions), or RLHF to refine validation mechanisms over time.
After successful validation, the intent execution phase applies the validated intent to the network infrastructure. The execution can be performed using API-based deployment for direct implementation, Open RAN network applications for network-specific functionalities, or negotiation-based execution via multiple LLM agents to resolve potential conflicts between different intents before deployment.
Figure 3: End-to-end GenAI framework for intent-driven network management.
III-A Hierarchical RAN Integration of the Proposed IDN Management Architecture
Modern hierarchical RAN architectures, such as Open RAN, use a multi-layered and modular design to support flexibility, scalability, and interoperability in next-generation wireless networks. These systems break down the traditional monolithic RAN into software-driven components that enable automation and AI-based optimization. In this work, we adopt a two-level architecture composed of a strategic controller and a tactical controller. The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations including handovers, resource allocation, and traffic steering. This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels, ensuring accurate interpretation and efficient execution of network intents.
We provide an example of implementing the three-step IDN management workflow using a custom hierarchical RAN architecture in Fig. 2. A human operator can observe the current network status or query for any relevant network information before providing an intent. The LLM module in the picture is given as an example. The operator may use logs, or another query-based database to gather information instead of using an LLM. After that, the operator can provide an intent that is received by the user interface application in the top-level Intelligent RAN Controller (IRC) named strategic controller. It then passes through the intent validation application. This application can be designed to perform intent validation via rule-based static methodologies or GenAI-based predictive analysis. Lastly, at the bottom, there is a tactical controller responsible for intent execution. In Fig. 2, we provide an example of intent execution via network applications in the tactical controller. Based on the intent, multiple network applications such as traffic steering, power allocation, and cell sleeping can be initiated and orchestrated to fulfill an intent.
IV A Case Study with Hierarchical Learning for IDN Management
Hierarchical application of GenAI algorithms can optimize complex decision-making by structuring learning at multiple levels. For instance, LLMs can interpret human intents, transformer-based predictors can validate network conditions, and decision transformers can execute intents. At the highest level, an LLM extracts key metrics and constraints from operator objectives. These insights guide a mid-tier forecasting model that predicts network behavior using historical and real-time data. At the lowest level, control algorithms generate policies for intent execution. This layered hierarchical approach improves efficiency, reduces computational overhead, and enables adaptive, real-time decisions
Fig. 3 presents an end-to-end GenAI framework for IDN management, which forms the foundation of our case study. Data collection follows a hierarchical reinforcement learning approach with decision-making split across two levels. The higher-level controller defines goals based on operator intents, such as increasing throughput or reducing power usage. The lower-level controller selects appropriate network applications depending on real-time conditions like traffic load, energy consumption, and packet loss. Five applications are used in this study: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover management. Each of them is developed using DRL. Traffic steering application distributes traffic intelligently across Radio Access Technologys (RATs) to maintain QoS, optimizing both delay and throughput. Cell sleeping reduces power consumption by deactivating underutilized base stations based on traffic load and queue length. Beamforming calculates optimal steering angles and selects matching array vectors using user location and the power allocation application improves total throughput through smart resource distribution. Lastly, the handover management application enhances energy efficiency by learning adaptive handover decisions. Once selected, each application independently adjusts its parameters using its own learning policy. More details can be found in [4].
Once an action is taken, the system evaluates its effectiveness by assessing whether the applied changes contribute to the desired network improvements. This evaluation is based on a reward system that provides feedback on whether the selected applications led to better performance. Over time, the system collects a large number of these decision-making instances to form a dataset that contains detailed records of past actions, network conditions, and resulting outcomes. The dataset is obtained via a custom-built simulation environment that mimics real-world network scenarios. Over time, this dataset can be extended and updated with real-world traces and trajectories to enhance its realism and applicability.
The collected data is structured into three distinct datasets. The first dataset is designed for fine-tuning an LLM and consists of query-response pairs and intent-based prompts to enable the model to accurately interpret operator-defined intents and provide relevant network insights. The second dataset comprises time series data, which includes key network parameters such as traffic load, power consumption, and packet loss percentage. This dataset is crucial for predictive intent validation, allowing the system to anticipate future network conditions and assess whether an intent can be successfully executed without degrading performance. The third dataset is a trajectory dataset used in the final stage, where control algorithms learn to optimize network application selection and orchestration. This dataset captures past decisions, network states, and corresponding outcomes to form the basis for training decision models that can make informed, goal-driven choices to achieve operator-defined objectives efficiently.
In this case study, we employ QLoRA to fine-tune an LLM (Llama 3.2 by Meta) to minimize computational resource requirements. Instead of traditional fine-tuning approaches that require high memory and processing power, QLoRA significantly reduces memory consumption without sacrificing performance. This fine-tuned model is further integrated with a RAG module  [^1] The RAG architecture is hybrid: it combines retrieval-based document fetching with the generative capabilities of the fine-tuned LLM., allowing it to access up-to-date network information and respond to dynamic queries with greater accuracy.
In the intent validation stage, the system evaluates whether the operator’s intent can be safely and effectively executed under current and predicted network conditions. This is done using a data-driven and predictive framework. First, the system continuously monitors key performance indicators including traffic load, packet loss percentage, and power consumption. These metrics are forecasted into future time slots using a transformer-based time series model that captures trends and relationships in temporal data. Once an intent is submitted, such as a request to improve throughput or reduce power usage, the system validates it against predicted network conditions. A multi-level check is performed where each intent is compared to predefined thresholds for the three metrics. For example, if a high throughput increase is requested during a period of already high traffic, the system may flag the intent as infeasible. A lookup table is used to cross-reference predicted values and QoS outcomes from past executions to enable fast and intelligent validation. This process helps ensure that only those intents are executed which are likely to maintain or improve overall network performance.
Finally, the last part of this case study includes the Mamba architecture for intent execution via network application orchestration. In particular, we use decision Mamba proposed in [13]. We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates decision Mamba as the high-level decision mechanism and a Decision Transformer (DT) as the low-level control transformer.
At a high level, decision Mamba serves as the goal-aware memory mechanism, responsible for retaining and retrieving critical past actions that have significantly contributed to achieving network optimization objectives. Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a Selective-State Space Modeling (SSSM) with a learnable dynamic memory retention mechanism. This allows the system to selectively remember only the most relevant past actions while discarding irrelevant ones. As a result, computational overhead gets reduced and decision efficiency improves.
At the low level, the decision transformer [14] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the decision Mamba. Once Mamba identifies the most impactful past action, it is passed to the control transformer, which uses a sequence modeling approach to predict the optimal action for the current state. By conditioning decisions on goal awareness rather than predefined reward sequences, the control transformer ensures that each selected action aligns with the operator’s high-level intent.
In this case study, when the operator provides an intent such as “Increase throughput by 10%”, the system sets this increased demand for throughput as its goal. It then recalls a past scenario with similar network conditions, such as high traffic load and moderate packet loss, where a specific combination of applications improved throughput. If the current network state does not support the intent, the system activates a feedback loop that either prompts the operator to revise the intent or uses the LLM to suggest feasible alternatives. The top-level GenAI block (decision Mamba) reviews the goal and the recent network context, retrieves the successful past action, and uses it to guide the present decision.
The control transformer described before has the current state of the network and the past action suggested by the top-level decision Mamba block. It uses this context to decide what to do next. It may decide to take the action of enabling a traffic steering application along with a power control application. This decision is tailored to current needs while being inspired by past success, aiming to fulfill the operator’s intent efficiently.
To evaluate the effectiveness of HDMGA, we compare it against two baseline approaches: Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation. The HDTGA framework, which is our prior method, utilizes a hierarchical decision transformer architecture. A meta-transformer searches past trajectories to retrieve the most relevant past action, which is then used to guide the control transformer in predicting future actions. The HRL with intent validation baseline follows a conventional HRL approach with a dedicated intent validation mechanism [4].
The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-RAT environment, serving $60$ users. The 5G New Radio (NR) operates at $3.5$ GHz (mid-band) and $30$ GHz (high-band) with bandwidths of $50$ MHz and $100$ MHz, respectively. The max transmission power of the 5G NR BS is $43$ dBm. On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $40$ MHz bandwidth and $38$ dBm maximum transmission power. Traffic types include video, gaming, voice, and vehicle-to-base station data traffic (Ultra-Reliable Low-Latency Communication (URLLC) use case), with packet inter-arrival times of $12.5$ ms, $40$ ms, $20$ ms, and $0.5$ ms, respectively, following Pareto, Uniform, and Poisson distributions.
In Fig. 4, we present spider plots of goal deviations for the proposed HDMGA and baseline methods. Delay-related goals (Fig. 4) include 2, 5, 7, and 11 ms. HDMGA consistently shows the lowest deviation, particularly excelling under stricter constraints. As delay targets relax (e.g., from 7 ms to 11 ms), all methods perform better due to increased flexibility, but HDMGA maintains superior accuracy. Similar trends are seen for energy-efficiency and throughput goals. Compared to HDTGA, HDMGA avoids attention overhead, reducing redundant computations. Additionally, Mamba’s linear-time processing enables efficient handling of long sequences, boosting throughput and making HDMGA more suitable for latency-sensitive applications.
Figure 4: Assessing the GenAI-based proposed three-step methodology in terms of delay goal deviations.
Figure 5: Performance comparison in terms of action inference time.
Fig. 5 presents a plot showing that the action inference time of the proposed method is lower than that of the baselines. The actions illustrated are not sequential decisions taken under identical network conditions. Instead, each action corresponds to an independent inference made under distinct network states at specific time points during the simulation. Since network conditions (e.g., traffic load, power consumption, packet loss) and operator intents vary over time, different methods naturally select varying combinations of applications based on context. This variability reflects each method’s adaptability to dynamic RAN environments. The figure demonstrates that HDMGA achieves faster decision-making, which is essential for near-real-time network automation. This efficiency is enabled by its memory-guided design, where decision Mamba selectively retains and retrieves only the most relevant past actions. This eliminates the need for exhaustive searches over large trajectory buffers. All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (10 ms to 1 s) [15].
A major challenge in systems with many network applications is the large and complex action space, which increases training time and hampers generalization. Our prior work addressed this using an attention-based HRL framework to filter irrelevant actions [4]. In this work, the proposed HDMGA further mitigates the issue through supervised learning approach on an offline-collected dataset, eliminating the need for online training during deployment.
V Conclusions
This paper introduces a hierarchical GenAI-driven framework for IDN management in 6G environments by integrating GenAI across intent processing, validation, and execution. By fine-tuning an LLM with QLoRA for efficient intent processing, using transformer-based predictors for validation, and deploying the HDMGA for execution, the proposed approach enhances network automation. The case study demonstrated HDMGA’s superior performance in achieving network objectives like delay reduction, increased energy efficiency, and throughput while reducing computational overhead. In our future works, we wish to integrate RLHF for refining intent validation and execution processes to enable real-time learning and adaptation based on operator feedback and network conditions.
Acknowledgment
This work has been supported by MITACS and Ericsson Canada, and NSERC Canada Research Chairs Program.

### v2

@@
-Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach
+Generative AI for Intent-Driven Network Management in 6G RAN: A Case Study on the Mamba Model
 Abstract
@@
 Intent-Driven Networks (IDNs) address this by translating high-level intents into optimization policies.
-Large Language Models (LLMs) can enhance this process by understanding complex human instructions to enable adaptive, intelligent automation.
+Large Language Models (LLMs) can enhance this process by understanding complex human instructions, enabling adaptive and intelligent automation.
 Given the rapid advancements in Generative AI (GenAI), a comprehensive survey of LLM-based IDN architectures in disaggregated Radio Access Network (RAN) environments is both timely and critical.
-This article provides such a survey, along with a case study on a hierarchical learning-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution.
-Unlike most existing approaches that apply GenAI in the form of LLMs for intent processing only, we propose a hierarchical framework that introduces GenAI across all three stages of IDN.
-To demonstrate the effectiveness of the proposed IDN management architecture, we present a case study based on the latest GenAI architecture named Mamba.
-The case study shows how the proposed GenAI-driven architecture enhances network performance through intelligent automation, surpassing the performance of the conventional IDN architectures.
-Index Terms:
-Intent processing, Intent validation, Intent execution, Generative AI, Hierarchical learning
+This article provides such a survey, along with a case study on a selective State-Space Model (SSM)-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution.
+For the first time in the literature, we propose a hierarchical framework built on Mamba-SSM that introduces GenAI across all stages of the IDN pipeline.
+We further present a case study demonstrating that the proposed Mamba architecture significantly improves network performance through intelligent automation, surpassing existing IDN approaches.
+In a multi-cell 5G/6G scenario, the proposed architecture reduces quality of service drift by up to $70\%$, improves throughput by up to $80$ Mbps, and lowers inference time to 60607070 ms, outperforming GenAI, reinforcement learning, and non-machine learning baselines.
 I Introduction
@@
 Furthermore, manual configurations are costly, prone to errors and not scalable [1].
-Intent-Driven Network (IDN) management is emerging as a solution to these challenges [1].
-An intent defines desired outcomes by specifying target metrics without detailing implementation steps to ensure both human comprehension and machine interpretability [4].
-Advances in Natural Language Processing (NLP) help operators express these intents more effectively.
-This reduces manual work, lowers errors, and speeds up configuration changes and service deployments [4, 5].
-Disaggregated Radio Access Networks (RANs), like Open RAN with its layered structure, are well-suited for IDN management.
-High-level controllers can interpret operator goals and check if the intent is achievable based on real-time conditions.
-Low-level controllers or functions then create the best network policies to meet these goals.
-This approach improves flexibility, scalability, and decision-making efficiency.
-On the other hand, Generative Artificial Intelligence (GenAI) has led to the rapid advancement of Large Language Models (LLMs) and revolutionized the field of NLP.
-As they have an outstanding ability to understand complex human instructions in natural language, IDN management is leaning towards LLMs instead of using fixed Service Level Agreement (SLA)-based approaches.
-For example, Dzeparoska et al.
-have introduced a pipeline that utilizes an LLM to process intents into structured and policy-based abstractions, which are next linked with Application Programming Interfaces (APIs) for execution [6].
-Also, in our previous work we have introduced an IDN management framework that uses LLM few-shot learning for intent processing from a human network operator [4].
-While these works use LLMs for intent processing, it can be highly beneficial to introduce GenAI not only as an intent processor via LLM but also we can take IDN management to the next level by introducing GenAI solutions in the other two stages namely intent validation and intent execution.
-Figure 1: Three-step methodology for intent-driven network management.
-With the inspiration for introducing efficient GenAI-based IDN management schemes via hierarchical disaggregated RAN environments, this work first reviews the state-of-the-art approaches that use GenAI for IDN management.
-Next, we introduce a three-step methodology for IDN management, encompassing intent processing, validation, and execution (see Fig.
-1 for details).
-After that, we present how these three key steps can be implemented using a hierarchical RAN architecture with the help of multiple GenAI algorithms.
-Lastly, we present a case study that shows the implementation of these steps.
-In particular, we process intents in a memory-efficient manner where an LLM is fine-tuned with a custom dataset using Quantized Low Rank Adaptation (QLoRA) [7].
-For the intent validation, a transformer-based time series predictor has been used in the case study which predicts crucial network parameters in future time slots and based on the predictions validates intents.
-Lastly, for intent execution, we generate network policies using Mamba, a GenAI architecture based on state space models for its demonstrated effectiveness compared to the transformer architecture in terms of memory and computation efficiency [8].
-The core contributions of this paper are as follows: it presents a precise, yet comprehensive survey of LLM-based IDN architectures, highlighting current limitations and identifying opportunities for integrating broader GenAI techniques across the stages of intent processing, validation, and execution.
-Furthermore, it introduces a three-step methodology for IDN management that encompasses these stages, integrated within a hierarchical RAN architecture.
-To validate the proposed framework, a case study demonstrating the utility of the proposed framework using decision Mamba has been presented for the first time in the literature.
+Intent-Driven Network (IDN) management is emerging as a solution to the challenges [1].
+An intent defines a desired network outcome by specifying target metrics without detailing the underlying implementation steps, thereby ensuring both human comprehensibility and machine interpretability [3].
+Advances in Natural Language Processing (NLP), driven primarily by the rise of Large Language Models (LLMs) have enabled operators to express intents more naturally and effectively.
+Beyond intent processing via LLMs, recent works have incorporated other forms of Generative Artificial Intelligence (GenAI) across the IDN pipeline, including transformer-based traffic prediction for intent validation [3] and policy execution via Decision Transformers [19].
+These GenAI-enabled components in IDN management collectively reduce manual workload, lower configuration-related errors, and accelerate service deployment and adaptation [4].
+However, an important limitation in existing works is the adoption of transformer-based GenAI models at every stage of IDN, including intent processing, predictive validation, and policy execution.
+This introduces a quadratic complexity bottleneck that increases inference latency and limits scalability as sequence lengths grow.
+To address this issue, a selective State Space Model (SSM) named Mamba is selected.
+Mamba provides a suitable alternative by enabling linear-time inference and efficient long-range dependency modeling.
+Mamba extends the SSM family by maintaining a recurrent hidden state while replacing fixed linear dynamics with a selective state-update mechanism.
+This design replaces the Transformer’s quadratic self-attention with a fast and selectively updated memory stream [9].
+In LLMs, this enables long-context reasoning without computational explosion and supports efficient processing of large text inputs, logs, and multi-page instructions.
+In decision-making settings, such as Decision Mamba for Reinforcement Learning (RL) [16], selective memory retains only historically meaningful transitions and avoids uniform weighting of past states.
+In time-series forecasting, Mamba captures both smooth temporal trends and sudden spikes without relying on positional encodings that limit extrapolation [20].
+Across these domains (LLM, time-series forecasting, and control decision making), Mamba addresses the structural limitations of Transformers while offering a faster, lighter, and more stable alternative.
+Motivated by the need for more efficient GenAI techniques in hierarchical and disaggregated Radio Access Network (RAN) environments, this work introduces Mamba as an alternative to transformer-based algorithms for IDN management.
+We begin with a brief review of state-of-the-art GenAI methods used across the IDN workflow.
+We then propose a three-stage Mamba-driven methodology that covers intent processing, validation, and execution.
+Building on this methodology, we also show how each stage can be realized within a hierarchical RAN architecture through multiple specialized Mamba-based models.
+A case study further demonstrates the practical realization of the entire pipeline.
+Specifically, intents are interpreted using a parameter-efficient Mamba-based LLM fine-tuned on a custom dataset [21].
+Intent feasibility is assessed using a Mamba-based time-series predictor that forecasts key network parameters and validates operator objectives accordingly.
+Finally, intent execution is performed via Decision Mamba, for its superior memory and computational efficiency compared to transformer-based approaches [9].
+The core contributions of this paper are threefold.
+First, it presents a precise, yet comprehensive survey of GenAI-based IDN architectures, highlighting their main features, limitations and applications.
+Along with the survey, we also present open research directions, practical deployment considerations, and standardization implications for GenAI-enabled IDN management, offering a roadmap for future development in this domain.
+Second, we introduce the first Mamba-driven IDN framework in the literature and demonstrate how it can replace traditional transformer-based components to deliver linear-time inference with improved long-range reasoning.
+Third, we show the effectiveness of the proposed framework through a detailed case study within a hierarchical RAN architecture that showcases the end-to-end workflow of Mamba-based IDN management and highlights its computational and operational advantages.
+Unlike prior GenAI-enabled IDN frameworks that rely on transformer-based models and typically introduce intelligence at only one or two stages of the IDN pipeline, this work is the first to systematically replace transformers across intent processing, validation, and execution using the Mamba models.
+This unified Mamba-driven architecture eliminates quadratic attention overhead, enables faster inference, and improves crucial Key Performance Indicators (KPIs).
+In our evaluation across a multi-cell Fifth-Generation (5G)/6G scenario, the proposed Mamba–enabled IDN architecture delivers substantial performance gains across all major KPIs.
+The framework consistently achieves the lowest Quality-of-Service (QoS) drift, reducing intent-violation rates by up to $70\%$ compared to the baselines.
+It further enhances network throughput by 45458080 Mbps under moderate to high traffic loads and achieves up to two times higher energy efficiency relative to the baselines.
+In addition, the proposed Mamba architecture for IDN management enables fast inference, lowering decision latency to 60607070 ms, which is considerably better than the baselines as well.
 II GenAI-Based Intent-driven Network Management Schemes
-Intent-based networking connects end-user service requests with network operations.
-Nijah et al.
-states that NLP-based methods offer the most flexibility to enable intent expression in natural language [2].
-LLMs have revolutionized NLP and they surpass traditional methods in flexibility and accuracy.
-Their ability to perform zero-shot, few-shot, and fine-tuned learning makes them state-of-the-art for tasks like intent recognition and processing.
-Existing research on IDN management schemes vastly uses LLMs to understand intents from the operators.
-While a few GenAI paradigms explore techniques like diffusion models and neuro-symbolic reasoning, the majority of current research in IDN management focuses on LLMs.
-Therefore, in this section, we review the LLM-centric approaches for IDN management in the telecommunication domain.
-Subsequently, we explore the potential application of LLM techniques or GenAI algorithms which to date, have not been widely explored in the literature.
-We also summarize the LLM-centric IDN management approaches, their main features, challenges, and applications in Table I.
+LLMs have transformed IDN management due to their ability to perform zero-shot, few-shot, and fine-tuned learning.
+They now represent the state-of-the-art for intent recognition and processing, and most existing IDN studies employ them to interpret operator intents.
+The literature has also evolved beyond intent processing, incorporating AI-based modules for validating and executing intents [1, 3].
+Recently, a full-scale GenAI-driven solution was proposed in [19].
+Motivated by these developments, this section first reviews works that focus specifically on GenAI-based methods directly supporting the IDN management workflow across intent processing, validation, and execution.
+We summarize reviewed works in terms of their main features, challenges, and applications in Table I.
+Note that we have selected works that have used GenAI in at least one of the stages of the IDN management workflow.
+Subsequently, in this section, we explore the open research directions, 3GPP and Open RAN standard implications, and practical deployment considerations for GenAI-based IDN management.
 TABLE I: LLM-Based IDN Management Approaches: Features, Challenges, and Applications
 Category Main Features Challenges Applications
-LLM-assisted intent processing and management for Fifth-Generation (5G) core networks [9, 10] Supports multiple intent transmission to the model.
+LLM-assisted intent processing and management for 5G core networks [11, 12] Supports multiple intent transmission to the model.
 Enhances network orchestration with semantic routing.
 Prevents LLM hallucination.
-Includes a RAG module.
-Reliance on LLMs which are not inherently designed for network control and automation.
+Includes a Retrieval Augmented Generation (RAG) module.
+Reliance on LLMs, which are not inherently designed for network control and automation.
 Distinguishing between similar or overlapping intents.
@@
 Mapping intents to edge-based service deployments.
-LLM-Driven multi-agent and negotiation-based intent management for 6G networks [11, 5] Employs MARL and LLM-driven negotiation.
+LLM-Driven multi-agent and negotiation-based intent management for 6G networks [13, 4] Employs Multi-Agent RL (MARL) and LLM-driven negotiation.
 Resolves conflicting intents in multi-stakeholder settings.
@@
 Autonomous resource management.
-End-to-end AI-based intent-driven network automation [1, 2, 4] Converts high-level intents into actionable network policies.
+End-to-end AI-based intent-driven network automation [1, 3, 19, 22] Converts high-level intents into actionable network policies.
 Verifies if an intent is feasible given the network’s current or predicted state.
@@
 Generative AI and multimodal intent-based network management [15] Uses multimodal generative AI for intent translation.
-LLMs are fine-tuned with few-shot learning to adapt to networking use cases.
+LLMs are tuned with few-shot learning.
 High dependency on predefined templates.
-Expansion to include multimodal inputs (e.g., images, topology files).
+Expansion to include multimodal inputs.
 Automation of the deployment and management of network slices.
-II-A State-of-the-art LLM-Based Approaches for IDN Management
+II-A State-of-the-art GenAI-Based Approaches for IDN Management
 II-A1 LLM-Assisted Intent Processing and Management for 5G Core Networks
@@
 A customized LLM model is designed to interpret and translate user intents into actionable network policies.
-By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks, such as deployment, modification, performance assurance, and feasibility checking.
+By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks.
 Semantic routing is introduced to refine LLM-assisted intent-based networking in [12].
@@
 To address these challenges, the proposed framework integrates a semantic router, which ensures deterministic decision-making by routing extracted intents through predefined pathways rather than relying solely on LLM-generated responses.
-This hybrid architecture improves reliability, reduces latency, and enhances accuracy in 5G core management and orchestration.
 II-A2 LLM-Driven Multi-Agent and Negotiation-Based Intent Management
@@
 Another study [4] introduces a comprehensive LLM-driven intent life cycle management system designed to handle all stages of intent processing, from decomposition and translation to negotiation, activation, and assurance.
-By shifting away from rigid JSON/YAML intent definitions, the system proposed in [5] enables natural language-driven network orchestration.
-Users can specify operational goals without requiring deep technical expertise.
 II-A3 End-to-End AI-Based Intent-Driven Network Automation
-Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), Multi-Agent RL (MARL), and Hierarchical RL (HRL) to realize fully automated network management.
-While processing intents is crucial, ensuring their feasibility and impact is equally important [4].
-A recent study introduces a transformer-based time series predictor to validate intents before execution.
-This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations (e.g., increasing energy efficiency, improving throughput) do not negatively affect service quality.
+Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), MARL, and Hierarchical RL (HRL) to realize fully automated network management.
+A recent study introduces a transformer-based time series predictor to validate intents before execution [3].
+This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations do not negatively affect service quality.
 Once an intent is validated, an HRL framework selects and triggers appropriate network optimization applications such as beamforming, traffic steering, and power control.
 An attention-based HRL model filters out suboptimal actions to reduce computational overhead while maximizing efficiency.
+Zhang et al.
+propose an intent-driven closed-loop control and management framework for 6G Open RAN, which integrates AI across all stages of the IDN pipeline [22].
+Intent translation is performed using a long short-term memory model with morphological rules, while intent verification and decomposition rely on AI-assisted reasoning via event calculus.
+The refined objectives feed into a control loop, where a Deep-Q-Network (DQN)-based network application (near-real-time control) executes slice-level resource allocation.
+Another recent contribution is proposed in [19], which demonstrates a fully GenAI-enabled workflow for intent processing, validation, and execution.
+The framework integrates a Quantized Low Rank Adaptation (QLoRA)-fine-tuned LLM for intent interpretation, an Informer model for long-horizon traffic prediction, and a goal-aware Decision Transformer (DT) for RAN control.
+By combining large-scale language understanding with predictive analytics and sequential decision-making, this work illustrates how GenAI can support seamless end-to-end automation in 6G Open RAN environments.
 Another breakthrough in end-to-end AI-driven network automation is the integration of multi-agent learning frameworks [1], where AI-driven network agents negotiate and resolve conflicting intents in real-time.
-This is particularly essential in multi-tenant 6G networks, where various stakeholders (e.g., operators, service providers, enterprise users) have competing resource demands.
-The study conducted in [1] proposes a multi-domain orchestration model, where LLM-driven agents representing RAN, core, and transport domains collaborate to optimize resource allocation and maintain Quality-of-Service (QoS).
+This is particularly essential in multi-tenant 6G networks, where various stakeholders have competing resource demands.
 II-A4 Generative AI and Multimodal Intent-Based Network Management
-A multimodal intent recognition framework presented in [12] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators to specify intents through text prompts, graphical designs, or configuration files.
+A multimodal intent recognition framework presented in [15] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators for specifying intents through text prompts, graphical designs, or configuration files.
 AI dynamically refines intents to ensure accuracy while aligning with service-level objectives.
@@
 The framework enables zero-touch deployment to map business requirements to network slice templates.
-II-B Potential GenAI Paradigms for Enhancing IDN Management
-The key limitations of the LLM-based IDN management schemes that we discussed in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements.
-To address these issues, we recommend some possible solutions from the GenAI perspective.
-$\bullet$ Retrieval Augmented Generation (RAG) and separate intent refinement/ validation scheme: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination.
-Furthermore, a multi-step intent refinement process can be introduced where a second model (e.g., a smaller LLM or rule-based checker) cross-checks intent translations before execution.
+II-B Open Research Directions for Enhancing IDN Management
+The key limitations of the GenAI-based IDN management schemes discussed so far in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements.
+To address these issues, we recommend some possible measures as open research directions.
+$\bullet$ RAG: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination.
+Particularly, there has not been any work in the IDN literature that has explored Agentic RAG (using LLM agent to choose between databases).
 $\bullet$ Reinforcement Learning with Human Feedback (RLHF): Since re-training is not always an option due to computational resource scarcity, RLHF can be a powerful technique for improving LLM-based IDN management to address the issue of real-time adaptation and feedback loops.
-We can define a reward function based on network Key Performance Indicators (KPIs) (e.g., latency, throughput, energy efficiency).
+We can define a reward function based on network KPIs (e.g., latency, throughput, energy efficiency).
 Then, a feedback loop can be implemented where operators and simulated environments score the LLM ’s intent outputs.
@@
 $\bullet$ LLM quantization and compression: Computation overhead caused by LLM deployments can be handled by implementing 4-bit quantization techniques (QLoRA, Generative Pre-trained Transformer-Quantization (GPTQ)) to reduce LLM model size while maintaining high accuracy.
-III Three-step Methodology for Intent-driven Network Management
-In this section of the paper, we present a three-step framework to introduce IDN management in modern-day communication systems.
-In particular, we propose a three-fold strategy consisting of intent processing, intent validation, and intent execution via network optimization policy generation.
+$\bullet$ Mamba models: Mamba models can be incorporated for reduced memory overhead and faster inference instead of the transformers in the IDN management pipeline.
+II-C Practical deployment considerations
+Practical deployment of GenAI-enabled IDN management must account for the strict latency requirements of RAN environments, where many control-loop decisions must be completed within the near-real-time window of 1010ms to $1$s.
+These constraints limit the size and complexity of deployable LLMs and predictive models.
+Another key challenge is ensuring robustness against QoS and intent drift, as real-world network conditions and operator objectives may shift over time.
+To maintain reliable operation, deployment pipelines must incorporate lightweight model updates, drift detection mechanisms, and safeguards that ensure GenAI-driven decisions remain compliant with QoS guarantees.
+II-D 3GPP and Open RAN standard implications
+GenAI-enabled IDN management has several implications for ongoing 3GPP and Open RAN standardization efforts.
+First, 3GPP Release 17 and beyond define intent-driven management services (TS 28.312) , but current specifications assume rule-based intent translation.
+Integrating LLM-based intent processing and SSM-driven validation requires extending these standards to support natural-language intent formats, semantic models, and AI-assisted intent interpretation pipelines.
+Second, predictive validation and goal-aware decision-making must align with 3GPP’s management data models, where AI-generated recommendations map to standardized network resource models and service exposure Application Programming Interfaces (APIs).
+From the Open RAN perspective, introducing GenAI into the intent execution loop affects the interfaces between the service management orchestrator, non-real-time intelligent controller, near-real-time intelligent controller, and network applications inhabiting in those controllers.
+Mamba-based predictors and decision models must comply with O-RAN open interfaces such as A1 (connects the non and near-real-time controllers), E2 (links the near-real-time controller with the distributed RAN nodes (O-RAN Distributed Unit (O-DU)/O-RAN Centralized Unit (O-CU)) for cross-layer orchestration.
+Moreover, near-real-time latency requirements (1010ms–11s) impose constraints on the deployment of GenAI models that motivate faster inference and hardware-aware optimization.
+III Three-step Methodology for IDN Management and its RAN Integration with GenAI
+In this section of the paper, we discuss a three-step framework consisting of intent processing, validation and execution to introduce IDN management and how GenAI can be integrated within it.
+This section will help the readers have the necessary background to understand the Mamba-based case study to be presented in the next section.
 Figure 1: Integration of the proposed IDN management framework with hierarchical RAN architecture.
@@
 $\bullet$ Intent execution: Intent execution is the final stage of IDN management, where validated intents are transformed into real-time network actions.
-This phase involves invoking appropriate network optimization policies to dynamically configure resources, and ensure that the desired service-level objectives are met efficiently.
-Intent execution can be realized through network function orchestration, AI-driven decision-making, or Reinforcement Learning (RL)-based adaptation.
-There are multiple ways to execute intents depending on the network architecture and management framework.
-The system can trigger predefined network applications such as traffic steering, beamforming, power control, and slicing management for intent execution.
-Also, LLMs can generate and refine network policies dynamically.
-Another approach for multi-operator environments can be to utilize RL-based agents that can negotiate optimal execution strategies to ensure efficient network optimization.
-To represent this three-step methodology, we presented Fig.1 in the Introduction, which illustrates the step-by-step process of intent-driven network management.
-The process begins with intent processing, where network intents can be expressed in natural language, structured formats (JSON, XML), or multi-modal data such as topology diagrams and configuration files.
-To process the intents, different methodologies can be applied, including LLMs for natural language understanding, rule-based approaches using predefined logic, graph-based methods for structured representation of dependencies, or finite state machines to model intent workflows as a sequence of state transitions.
-Once an intent is processed, it moves to intent validation which can be carried out using rule-based validation (checking against static policies), constraint-based validation (ensuring resource availability and compliance), predictive validation (using AI/ML models to anticipate network conditions), or RLHF to refine validation mechanisms over time.
-After successful validation, the intent execution phase applies the validated intent to the network infrastructure.
-The execution can be performed using API-based deployment for direct implementation, Open RAN network applications for network-specific functionalities, or negotiation-based execution via multiple LLM agents to resolve potential conflicts between different intents before deployment.
+This phase involves invoking appropriate network optimization policies to dynamically configure resources and ensure that the desired service-level objectives are met efficiently.
+Intent execution can be realized through network function orchestration, AI-driven decision-making, or RL-based adaptation.
 Figure 2: End-to-end GenAI framework for intent-driven network management.
-III-A Hierarchical RAN Integration of the Proposed IDN Management Architecture
+III-A Hierarchical RAN Integration with the three-step IDN Management Architecture
 Modern hierarchical RAN architectures, such as Open RAN, use a multi-layered and modular design to support flexibility, scalability, and interoperability in next-generation wireless networks.
@@
 In this work, we adopt a two-level architecture composed of a strategic controller and a tactical controller.
-The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations including handovers, resource allocation, and traffic steering.
-This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels, ensuring accurate interpretation and efficient execution of network intents.
+The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations, including handovers, resource allocation, and traffic steering.
+This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels.
 We provide an example of implementing the three-step IDN management workflow using a custom hierarchical RAN architecture in Fig.
-2.
+1.
 A human operator can observe the current network status or query for any relevant network information before providing an intent.
 The LLM module in the picture is given as an example.
-The operator may use logs, or another query-based database to gather information instead of using an LLM.
+The operator may use logs or another query-based database to gather information instead of using an LLM.
 After that, the operator can provide an intent that is received by the user interface application in the top-level Intelligent RAN Controller (IRC) named strategic controller.
@@
 In Fig.
-2, we provide an example of intent execution via network applications in the tactical controller.
+1, we provide an example of intent execution via network applications in the tactical controller.
 Based on the intent, multiple network applications such as traffic steering, power allocation, and cell sleeping can be initiated and orchestrated to fulfill an intent.
-IV A Case Study with Hierarchical Learning for IDN Management
-Hierarchical application of GenAI algorithms can optimize complex decision-making by structuring learning at multiple levels.
-For instance, LLMs can interpret human intents, transformer-based predictors can validate network conditions, and decision transformers can execute intents.
-At the highest level, an LLM extracts key metrics and constraints from operator objectives.
-These insights guide a mid-tier forecasting model that predicts network behavior using historical and real-time data.
-At the lowest level, control algorithms generate policies for intent execution.
-This layered hierarchical approach improves efficiency, reduces computational overhead, and enables adaptive, real-time decisions
+IV A Case Study on Mamba Model for IDN Management
+Hierarchical use of GenAI algorithms can simplify complex decision-making by splitting learning across multiple layers.
+At the top layer, LLMs interpret human intents and extract key metrics and constraints.
+These insights guide a mid-tier time-series predictor that analyzes historical and real-time data to validate network conditions.
+At the lowest layer, control algorithms translate validated intents into executable policies.
+This layered design improves efficiency, lowers computational overhead, and supports adaptive, real-time decision-making.
 Fig.
-3 presents an end-to-end GenAI framework for IDN management, which forms the foundation of our case study.
-Data collection follows a hierarchical reinforcement learning approach with decision-making split across two levels.
-The higher-level controller defines goals based on operator intents, such as increasing throughput or reducing power usage.
-The lower-level controller selects appropriate network applications depending on real-time conditions like traffic load, energy consumption, and packet loss.
-Five applications are used in this study: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover management.
-Each of them is developed using DRL.
-Traffic steering application distributes traffic intelligently across Radio Access Technologys (RATs) to maintain QoS, optimizing both delay and throughput.
-Cell sleeping reduces power consumption by deactivating underutilized base stations based on traffic load and queue length.
-Beamforming calculates optimal steering angles and selects matching array vectors using user location and the power allocation application improves total throughput through smart resource distribution.
-Lastly, the handover management application enhances energy efficiency by learning adaptive handover decisions.
-Once selected, each application independently adjusts its parameters using its own learning policy.
-More details can be found in [4].
-Once an action is taken, the system evaluates its effectiveness by assessing whether the applied changes contribute to the desired network improvements.
-This evaluation is based on a reward system that provides feedback on whether the selected applications led to better performance.
-Over time, the system collects a large number of these decision-making instances to form a dataset that contains detailed records of past actions, network conditions, and resulting outcomes.
-The dataset is obtained via a custom-built simulation environment that mimics real-world network scenarios.
-Over time, this dataset can be extended and updated with real-world traces and trajectories to enhance its realism and applicability.
-The collected data is structured into three distinct datasets.
-The first dataset is designed for fine-tuning an LLM and consists of query-response pairs and intent-based prompts to enable the model to accurately interpret operator-defined intents and provide relevant network insights.
-The second dataset comprises time series data, which includes key network parameters such as traffic load, power consumption, and packet loss percentage.
-This dataset is crucial for predictive intent validation, allowing the system to anticipate future network conditions and assess whether an intent can be successfully executed without degrading performance.
-The third dataset is a trajectory dataset used in the final stage, where control algorithms learn to optimize network application selection and orchestration.
-This dataset captures past decisions, network states, and corresponding outcomes to form the basis for training decision models that can make informed, goal-driven choices to achieve operator-defined objectives efficiently.
-In this case study, we employ QLoRA to fine-tune an LLM (Llama 3.2 by Meta) to minimize computational resource requirements.
-Instead of traditional fine-tuning approaches that require high memory and processing power, QLoRA significantly reduces memory consumption without sacrificing performance.
-This fine-tuned model is further integrated with a RAG module  [^1] The RAG architecture is hybrid: it combines retrieval-based document fetching with the generative capabilities of the fine-tuned LLM., allowing it to access up-to-date network information and respond to dynamic queries with greater accuracy.
-In the intent validation stage, the system evaluates whether the operator’s intent can be safely and effectively executed under current and predicted network conditions.
-This is done using a data-driven and predictive framework.
-First, the system continuously monitors key performance indicators including traffic load, packet loss percentage, and power consumption.
-These metrics are forecasted into future time slots using a transformer-based time series model that captures trends and relationships in temporal data.
-Once an intent is submitted, such as a request to improve throughput or reduce power usage, the system validates it against predicted network conditions.
-A multi-level check is performed where each intent is compared to predefined thresholds for the three metrics.
-For example, if a high throughput increase is requested during a period of already high traffic, the system may flag the intent as infeasible.
-A lookup table is used to cross-reference predicted values and QoS outcomes from past executions to enable fast and intelligent validation.
-This process helps ensure that only those intents are executed which are likely to maintain or improve overall network performance.
+2 shows the end-to-end GenAI framework for IDN management used in our case study.
+Data collection follows an HRL structure with two decision-making levels.
+The higher-level controller sets goals based on operator intents, such as raising throughput or lowering power use.
+The lower-level controller chooses the appropriate network application based on traffic load, energy efficiency, or packet loss.
+Five applications are used: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover.
+Each is implemented using DRL.
+Traffic steering distributes traffic across different radio access technologies to improve delay and throughput.
+Cell sleeping reduces power by disabling underutilized base stations.
+Beamforming computes optimal angles using user location, while power allocation improves total throughput through smart resource assignment.
+The handover management application enhances energy efficiency through adaptive handover decisions.
+After selection, each application independently adjusts its parameters using its own policy.
+More details are available in [3].
+After each action, the system evaluates its impact by checking whether it improves the targeted network metrics.
+A reward mechanism measures how effective the selected applications were in enhancing performance.
+Over time, this process produces a large dataset containing past actions, network states, and observed outcomes.
+The dataset is generated using a custom simulator designed to mimic realistic network conditions and can later be enriched with real-world traces to further improve accuracy.
+The collected data is organized into three datasets.
+The first supports LLM fine-tuning and contains query–response pairs and intent-driven prompts to help the model interpret operator goals.
+The second dataset contains time-series measurements of parameters such as traffic load, packet loss, and power consumption.
+This data enables predictive intent validation by forecasting whether an intent is safe to execute.
+The third dataset is a trajectory set that trains the control algorithms.
+It records past decisions, states, and outcomes, enabling decision models to learn effective application selection that meets operator objectives.
+In this case study, we adopt Mamba-Parameter Efficient Fine-Tuning (PEFT) to fine-tune a pretrained Mamba-based LLM for intent-driven RAN management [21].
+Instead of updating all model parameters, which would require high memory and computing, Mamba-PEFT modifies only selected components inside the Mamba block.
+These include the input and output projections, specific state-space transitions, or a small number of additional state dimensions.
+By limiting the number of trainable parameters, the method reduces memory usage and keeps the fine-tuning process efficient while preserving the model’s long-context reasoning ability.
+Since Mamba processes sequences in linear time, the overall adaptation becomes more scalable than transformer-based alternatives.
+The fine-tuned model is integrated with a RAG module, which provides access to current network information, configurations, and logs.
+With both selective parameter adaptation and retrieval-augmented knowledge, the resulting Mamba-PEFT model delivers accurate and context-aware responses to operator queries while remaining suitable for use in hierarchical RAN environments.
+In the intent validation stage, the system evaluates whether an operator’s intent can be executed safely under both current and predicted network conditions.
+KPIs such as traffic load, packet loss, and power consumption are continuously monitored and forecasted using Mamba4Cast [20].
+Mamba4Cast is a Mamba-based zero-shot time-series foundation model built on selective state-space representations.
+It combines causal convolution layers with lightweight Mamba-based encoder blocks to capture multi-scale temporal patterns and generates multi-horizon forecasts in a single pass.
+When an operator submits an intent to increase throughput or reduce power usage, the predicted KPI trajectories are checked against predefined feasibility thresholds.
+If the forecasts indicate a risk of performance degradation, such as requesting higher throughput during a period of high expected load, the system marks the intent as infeasible.
+A lookup table that maps predicted KPI patterns to known quality-of-service outcomes from past executions further accelerates the decision.
+Through this combination of Mamba-based time-series forecasting and historical pattern matching, the validator ensures that only intents suitable for future conditions proceed to execution.
 Finally, the last part of this case study includes the Mamba architecture for intent execution via network application orchestration.
-In particular, we use decision Mamba proposed in [13].
-We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates decision Mamba as the high-level decision mechanism and a Decision Transformer (DT) as the low-level control transformer.
+In particular, we use the Decision Mamba proposed in [16].
+We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates Decision Mamba as the high-level decision mechanism and a DT as the low-level control transformer.
 At a high level, Decision Mamba serves as the goal-aware memory mechanism, responsible for retaining and retrieving critical past actions that have significantly contributed to achieving network optimization objectives.
-Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a Selective-State Space Modeling (SSSM) with a learnable dynamic memory retention mechanism.
+Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a selective SSM with a learnable dynamic memory retention mechanism.
 This allows the system to selectively remember only the most relevant past actions while discarding irrelevant ones.
-As a result, computational overhead gets reduced and decision efficiency improves.
-At the low level, the decision transformer [14] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the decision Mamba.
+As a result, computational overhead gets reduced, and decision efficiency improves.
+At the low level, the DT [19] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the Decision Mamba.
 Once Mamba identifies the most impactful past action, it is passed to the control transformer, which uses a sequence modeling approach to predict the optimal action for the current state.
 By conditioning decisions on goal awareness rather than predefined reward sequences, the control transformer ensures that each selected action aligns with the operator’s high-level intent.
+The intent format, inputs, and outputs used across the pipeline are as follows.
+For intent processing, the input is either a natural-language query requesting contextual network information or an optimization intent specifying a KPI and a desired magnitude.
+The output is either an LLM-generated informational response or a structured representation identifying the target KPI and magnitude, respectively.
+For intent validation, the input is the parsed optimization intent together with predicted network KPIs generated by the forecasting module, and the output is a feasibility decision indicating whether the intent can be safely executed under current or predicted conditions.
+For intent execution, the input is the desired magnitude of improvement of a KPI, and the output is a set of executable network actions, including the selected applications and their control parameters, deployed in the RAN controllers.
 In this case study, when the operator provides an intent such as “Increase throughput by 10%”, the system sets this increased demand for throughput as its goal.
@@
 This decision is tailored to current needs while being inspired by past success, aiming to fulfill the operator’s intent efficiently.
-To evaluate the effectiveness of HDMGA, we compare it against two baseline approaches: Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation.
+Figure 3: Performance comparison of the proposed method against the baselines in terms of goal deviation (network delay and energy efficiency).
+Figure 4: Performance comparison of the proposed method against the baselines in terms of throughput and reliability (QoS drift).
+To evaluate the effectiveness of HDMGA, we compare it against four baseline approaches.
+First of all, we adopt the deterministic rule-based method, which performs threshold-driven application triggering without any learning capability.
+This is referred to as Non-ML baseline in our performance graphs.
+Furthermore, we consider another baseline named StARformer where a transformer model is merged into an RL agent that predicts actions directly from recent state–action–reward tuples [23].
+We address this as Transformer-RL (TRL) in our performance graphs.
+The last two baselines are Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation.
 The HDTGA framework, which is our prior method, utilizes a hierarchical decision transformer architecture.
@@
 The HRL with intent validation baseline follows a conventional HRL approach with a dedicated intent validation mechanism [3].
-The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-RAT environment, serving $60$ users.
+The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-Radio Access Technology (RAT) environment, serving $60$ users.
 The 5G New Radio (NR) operates at $3.5$ GHz (mid-band) and $30$ GHz (high-band) with bandwidths of $50$ MHz and $100$ MHz, respectively.
 The max transmission power of the 5G NR BS is $43$ dBm.
-On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $40$ MHz bandwidth and $38$ dBm maximum transmission power.
+On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $20$ MHz bandwidth and $38$ dBm maximum transmission power.
+The simulator models the behavior of the RAN using a 3GPP-aligned design that includes realistic traffic patterns, mobility, and Radio Frequency (RF) impairments.
+Four traffic types are generated each having its own inter-arrival time and distribution, which creates a mix of bursty, periodic, and low-latency traffic.
 Traffic types include video, gaming, voice, and vehicle-to-base station data traffic (Ultra-Reliable Low-Latency Communication (URLLC) use case), with packet inter-arrival times of $12.5$ ms, $40$ ms, $20$ ms, and $0.5$ ms, respectively, following Pareto, Uniform, and Poisson distributions.
-In Fig.
-4, we present spider plots of goal deviations for the proposed HDMGA and baseline methods.
-Delay-related goals (Fig.
-4) include 2, 5, 7, and 11 ms.
-HDMGA consistently shows the lowest deviation, particularly excelling under stricter constraints.
-As delay targets relax (e.g., from 7 ms to 11 ms), all methods perform better due to increased flexibility, but HDMGA maintains superior accuracy.
-Similar trends are seen for energy-efficiency and throughput goals.
-Compared to HDTGA, HDMGA avoids attention overhead, reducing redundant computations.
-Additionally, Mamba’s linear-time processing enables efficient handling of long sequences, boosting throughput and making HDMGA more suitable for latency-sensitive applications.
-Figure 4: Assessing the GenAI-based proposed three-step methodology in terms of delay goal deviations.
+User mobility follows 3GPP system-level guidelines, where User Equipment (UE) positions are updated every Transmission Time Interval (TTI) and movement directly affects path loss, Doppler shift, beam directions, and handover events.
+The RF environment includes Clustered Delay Line (CDL)-based multipath fading, shadowing, thermal noise, beam misalignment, and multi-cell interference.
+These impairments evolve with mobility and produce realistic fluctuations in signal-to-interference-plus-noise ratio, throughput, and delay.
+Together, the traffic models, mobility behavior, and RF effects create a reliable and reproducible simulation environment for evaluating intent-driven RAN management.
+Our proof-of-concept is implemented using a multi-RAT MATLAB simulation (5G, LTE and CDL fading) integrated with Python-based Decision Mamba and Mamba-LLM modules.
+The LLM is fine-tuned using a 1.11.1k-sample dataset [19] of dynamic network queries, semi-static operational knowledge, and structured intents, while a Mamba-based predictor validates intent feasibility by forecasting load, power, and packet-drop trends.
+As presented in Fig.
+3, it can be observed that across both delay and energy-efficiency targets, the proposed HDMGA achieves the smallest goal deviations among all compared methods.
+This improvement is due to HDMGA ’s selective state-space modeling and goal-conditioned hierarchical design, which allow the agent to retain only the most relevant historical transitions and choose application combinations that directly contribute to meeting the operator’s intent.
+HDTGA performs moderately well but still shows higher deviations because its transformer-based attention weighs all past states uniformly and introduces unnecessary computational overhead.
+TRL experiences larger deviations since it lacks hierarchy, sub-goal decomposition, and selective memory, making it less effective when coordinating multiple network applications.
+HRL shows even higher deviations because it must explore the entire combinatorial action space without attention-based filtering or intent validation, which often results in suboptimal application choices under dynamic traffic conditions.
+The Non-ML heuristic baseline records the largest deviations for both metrics because its static threshold-based rules cannot adapt to changing network states or operator intent magnitudes.
+While only delay and energy-efficiency figures are presented, the same performance trend is consistently observed for throughput goal deviations as well.
+The QoS drift rate represents how much the network’s performance deviates from the level expected by the operator when an intent is executed.
+The percentage is computed by averaging the deviation over a monitoring window that spans several TTIs.
+Fig.
+4 compares the drift rates of all methods across different traffic loads.
+The proposed HDMGA maintains the lowest drift rates at every load level, indicating that it consistently meets the operator’s expected performance targets with minimal deviation.
+This reliability comes from its selective state-space modeling and goal-aware hierarchical design, which allow the agent to prioritize only the most relevant past transitions and choose the correct combination of applications even when the network becomes highly loaded.
+The throughput results in Fig.
+4 follow the same overall trend observed in the drift analysis.
+As the traffic load increases, HDMGA consistently achieves the highest throughput among all methods, reflecting its ability to select the most effective combination of RAN applications while respecting operator intent.
+Its selective memory mechanism helps preserve the most relevant context for scheduling and resource allocation, which leads to more efficient utilization of network capacity.
 Figure 5: Performance comparison in terms of action inference time.
@@
 This efficiency is enabled by its memory-guided design, where Decision Mamba selectively retains and retrieves only the most relevant past actions.
-This eliminates the need for exhaustive searches over large trajectory buffers.
-All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (10 ms to 1 s) [15].
-A major challenge in systems with many network applications is the large and complex action space, which increases training time and hampers generalization.
-Our prior work addressed this using an attention-based HRL framework to filter irrelevant actions [4].
-In this work, the proposed HDMGA further mitigates the issue through supervised learning approach on an offline-collected dataset, eliminating the need for online training during deployment.
+All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (1010ms to $1$s).
 V Conclusions
-This paper introduces a hierarchical GenAI-driven framework for IDN management in 6G environments by integrating GenAI across intent processing, validation, and execution.
-By fine-tuning an LLM with QLoRA for efficient intent processing, using transformer-based predictors for validation, and deploying the HDMGA for execution, the proposed approach enhances network automation.
-The case study demonstrated HDMGA’s superior performance in achieving network objectives like delay reduction, increased energy efficiency, and throughput while reducing computational overhead.
-In our future works, we wish to integrate RLHF for refining intent validation and execution processes to enable real-time learning and adaptation based on operator feedback and network conditions.
+This article presented a GenAI-driven framework for IDN management that integrates selective state-space modeling across intent processing, validation, and execution for the first time in the literature.
+By utilizing a fine-tuned Mamba-based LLM for intent interpretation, a Mamba-enabled predictor for proactive validation, and a Mamba-enhanced intent execution module, the proposed architecture delivers a unified and scalable approach to intelligent automation in 6G RAN environments.
+The case study demonstrated that this Mamba’s SSM-driven design significantly improves delay, throughput, energy efficiency, and reliability compared with existing IDN management approaches, while reducing computational overhead through efficient sequence modeling.
+Looking ahead, scaling the proposed architecture to larger deployments or conducting prototype-level validation (e.g., on Open RAN compliant testbeds) represents an important direction for future work.
 Acknowledgment
-This work has been supported by MITACS and Ericsson Canada, and NSERC Canada Research Chairs Program.
+This work has been supported by MITACS, Ericsson Canada, and NSERC Canada Research Chairs Program.